# Understanding Prompt Tuning and In-Context Learning via Meta-Learning

### v1

@@ Abstract
Keywords: prompting, prefix-tuning, in-context learning, meta-learning, Bayes, frontier models
@@ 1 Introduction
(a) Top: Performance of different tuning methods on Transformers, measured as excess log loss, i.e., regret (Eq. 10, lower is better). See bar plots for color legend. Top-right: Detailed Transformer results for last step within the tuning sequence length ${N_{\text{tune}}}$ and the last evaluation step ${N_{\text{eval}}}$. Right: Like above but for LSTM.
@@
Figure 1: Models pretrained on sequences from coins with uniform random bias (length ${N_{\text{train}}=100}$) are fine-tuned to the target task of a single coin with bias $0.2$ (tuning sequence length ${N_{\text{tune}}=50}$). Plots show prediction performance for tuned Transformers and LSTMs on the target task for different prefix- and weight-tuning methods. For both network types Soft Prompting (‘SoftPT’) is the only prefix-tuning method that leads to optimal performance. This holds up to the tuning sequence length, and degradations are only minor up to ${N_{\text{eval}}=200}$ steps. The corresponding soft prefix of length$=6$ outperforms even the best hard token prefix of the same length (‘HardPF’). Most weight-tuning methods also perform very well in this setting. See Section 4 for details on methods. Thick lines and bars show the median over $10$ tuning repetitions, thin lines individual repetitions, and shaded areas/bars show $25\%,75\%$ quantiles. See Fig. 2 for a visualization of models’ internal dynamics. Regret curves for the LSTM, similar to top-left panel, are shown in Fig. A5.
In this paper, we investigate prefix-tuning, where a short prompt prefix is tuned to maximize prediction accuracy over the subsequent tokens. We consider prefix search over hard tokens, and three soft prefix-tuning methods, including ‘Soft Prompting’, (Lester et al., 2021). The code to reproduce all our experiments is available at: https://github.com/google-deepmind/thunnini.
Our main contributions are:
  • We discuss how prompting can be understood as steering a Bayesian sequential predictor via its pretrained in-context adaptation mechanism that arises from meta-training, see Section 2.
  • We analyze theoretical conditions for the relationship between pretraining distribution and target task under which optimal prompting is and is not possible, see Section 3.
  • We confirm these theoretical conditions empirically via a series of educational experiments, and show that in the negative case weight-based fine-tuning can succeed. See Section 4.2.
  • We investigate mechanistic aspects of prompting LSTMs and Transformers, with experiments on both pretrained and untrained networks. In all cases, Soft Prompting can be much more effective than any hard token sequence of equal length, particularly for Transformers, where Soft Prompts can even cause untrained networks to behave as well-performing sequential predictors. See Section 4.2.
@@
As with any conceptual principle, the Bayesian view cannot exhaustively describe all in-context learning phenomena at (frontier model) scale: limited data and model expressivity, suboptimal optimization, out-of-distribution generalization, and other mechanistic aspects will have additional impact. Having said that, there are a number of recent investigations at LLM scale supporting the plausibility of the Bayesian view, such as Chan et al. (2022a) who argue that distributional properties of natural language can easily give rise to a meta-learning setting. The simplicity of our experimental setup allows to carefully control the data statistics, and compare neural predictors against exact Bayes-optimal predictors—both of which is not true for LLM-scale experiments. Additionally, it allows us to focus on fundamental aspects of prompting that arise even in idealized settings, and thus hold at any scale. We believe that this lays important fundamental groundwork, which will help future research to isolate and more effectively investigate additional, non-idealized, aspects of in-context learning and prompt tuning at frontier scale. We do not propose improved prompt tuning methods in this work since our aim is to study and understand existing ones.
@@ 2 Background: Memory-Based Meta-Learning
In this section we review how memory-based meta learning leads to Bayesian sequential predictors. Under this view, the role of prompt tuning is to facilitate the inference process of the target task.
@@
Let $\pi_{\theta}$ be a parametric sequential predictor, such as a neural network, which is a function with parameters $\theta$ that takes in an arbitrary-length sequence of $D$-dimensional float vectors and outputs a discrete distribution over the next (one-hot) token: $\pi_{\theta}:\{\mathbb{R}^{D}\}^{*}\rightarrow\Delta\mathcal{A}$ Typically, ${D=|\mathcal{A}|}$, meaning that one-hot tokens can be fed directly into the predictor. The predictor’s conditional distribution over the next token given a context $x_{<n}$ is given by evaluating the predictor with the given context (i.e., preforming a forward pass): ${P_{\theta}(x_{n}|x_{<n})=\pi_{\theta}(x_{<n})}$. Keep in mind that a neural net accepts not only hard tokens as inputs, but sequences of arbitrary vectors in $\mathbb{R}^{D}$ (‘soft tokens’), and will produce a distribution over the next hard token in either case.
@@
which is zero iff $\pi_{\theta}=\xi$. Any predictor that fulfills this is Bayes-optimal for $\xi$, and Eq. 1 and Eq. 2 provide a recipe for constructing an explicit Bayesian predictor, which is a mixture over one predictor per task, weighted by the posterior probability over the task given the context so far: ${P(\tau|x_{<n})\propto P(x_{<n}|\tau)P(\tau)}$. In many cases this recipe is analytically or computationally intractable. Memory-based meta-learning provides an alternative for obtaining an approximate Bayesian predictor simply through log loss minimization in a meta-learning loop:
@@
meaning the network’s prediction over the next token, given context $x_{<n}$, is (nearly) indistinguishable from an explicit Bayesian predictor. The meta-trained neural network thus implements a Bayes-optimal adaptive prediction algorithm via its activations only—without weight updates (in-context learning). Previous works have empirically verified that meta-trained LSTMs and Transformers can indeed reach Bayes-optimality through meta-training, e.g., Mikulik et al. (2020) for sequential prediction and decision-making (not covered in this paper, but the theory extends straightforwardly to loss functions other than log loss), Genewein et al. (2023) for piecewise stationary data sources (where models additionally have to infer task boundaries), and Grau-Moya et al. (2024) for variable-order Markov processes and for distributions over outputs of short programs for a simple universal Turing machine.
@@
Frontier models are implicitly meta-trained on samples from an unknown, rich and complex distribution of data generators (Chan et al., 2022a). The main concerns w.r.t. the applicability of the Bayesian viewpoint is that due to limited expressivity, limited data, suboptimal optimization, and off-distribution inputs, models may not converge to or operate in the Bayesian regime. While we strongly recommend carefully investigating these issues, the theory tells us that as models get better and better, they will get closer and closer to the Bayesian ideal, making it an important fundamental computational mechanism to understand.
@@ 3 Prompt Optimization: prefix-tuning
[…]
In this paper we use:
@@
that is, we are optimizing for a single target task that had support under the pretraining distribution. In this case, there always exists a sequence of hard tokens that causes the Bayesian posterior to concentrate sufficiently (in the limit a delta, see Appendix E ) for optimal prediction after the prefix. For sufficiently large $L$,
@@
The prefix can be found by performing the minimization in Eq. 7. Proof sketch: If the target distribution is a delta over one of the pretraining tasks (condition in Eq. 8), then the argument that is being minimized in Eq. 7 is the r.h.s. of Eq. 9. The minimum is obtained when Eq. 9 becomes an equality, which the case when the posterior ${\xi^{\text{Pre}}(\cdot|s_{1:L})}$ collapses to $\tau^{\text{Target}}$. This also implies ${D_{\text{KL}}(\xi^{\text{Target}}||\xi^{\text{Pre}}(\cdot|s_{1:L}))\approx 0}$ which is only possible iff $\xi^{\text{Target}}\approx\xi^{\text{Pre}}(\cdot|s_{1:L})$.
If the condition in Eq. 8 does not hold, optimal prompting may still be possible, but this strongly depends on the relationship between pretraining and target distribution and the model class. See Appendix E for an analysis of the Beta-Bernoulli case and beyond, including limits of prompting universal predictors. It is also possible to formulate two general theoretical negative cases:
@@
In the limit posteriors can often not remain (or become) multimodal. If the target distribution is multimodal, such as a finite target mixture over tasks, optimal prefix-tuning may not be possible, even if all mixture components have support under the pretraining distribution. For instance, if the prior ${P^{\text{Pre}}(\tau)}$ is log-concave and the likelihood function is log-concave, then the posterior is also log-concave, and thus unimodal. See Section 4.2 for an empirical demonstration. More generally, if the posterior collapses to a Dirac delta in the limit (as it does for a Beta-Bernoulli model, and is very likely to do in general if prompts are typical sequences—see Appendix E), the only target task distributions that are optimally promptable are deltas over a single pretraining task.
@@
In general, for a prefix prompt to be optimal, the model’s internal state after consuming the prefix needs to be a sufficient statistic for the target distribution, without causing subsequent internal dynamics to diverge. Pretraining determines a model’s state-update function, and thus imposes strong constraints w.r.t. possible manipulations via (hard token) inputs. These constraints can be partly overcome by using soft prefixes instead of hard tokens. As our results show, these carefully tuned off-distribution inputs that lead to off-distribution internal states, can be used to very effectively steer pretrained, and even untrained neural predictors. The limits of this mechanistic aspect, outside the conceptual Bayesian theory, are currently unclear—it could be that soft prefixes can very flexibly “reprogram” a pretrained model to arbitrary target distributions. Empirically, we find this not to be the case: while Soft Prompting does consistently improve prediction performance on the target task, it is still bound by the theoretical limitations pointed out in Appendix E: optimally adapting a predictor pretrained over uniform random coins to a mixture of two coins is not possible. How large the potential gains from Soft Prompting or other soft prefix-tuning methods can be is an empirical question. Finally, note that weight-based fine-tuning methods are able to modify the pretrained state-update and -readout mechanisms, which allows for more flexibility w.r.t. adapting a network to target distributions (see Section 4.2 for empirical demonstrations).
@@ 4 Experiments on Coin-Flip Sequences
We conduct a series of experiments where a neural network is first meta-trained over a pretraining distribution (Section 2) of coin flip sequences of length $N_{\text{pre}}=100$, and then prompt-tuned (Eq. 7) or weight-tuned (mini-batch based log-loss minimization) to a target distribution of coin flip sequences of length $N_{\text{tune}}=50$. After tuning, we evaluate tuned models on $2048$ sequences of length $N_{\text{Eval}}=200$ from the target distribution. Choosing ${N_{\text{Eval}}>N_{\text{tune}}}$ also allows to study how the solutions of different tuning methods generalize beyond the tuning length. Across experiments we use three different data distributions, two neural architectures, and nine tuning methods, which we now describe.
@@ 4.1 Experimental setup
We use coin-flip sequences $P(x_{1:N}|\tau)=\text{Bernoulli}(\tau)$ with three different distributions $P(\tau)$ throughout our experiments. Random coins: $P{(\tau)=\text{Beta}(1,1)}$, leading to a uniform distribution over coin biases. This is our pretraining distribution. The exact Bayesian predictor in this case is the Laplace predictor. Single coin: A single coin with bias $0.2$. This target distribution fulfills the condition in Eq. 8. Two-coin mixture: A mixture of two coins, one with bias $0.2$ and one with bias $0.8$, with equal mixing weights of $1/2$ each. This target distribution violates the condition in Eq. 8. All tasks have binary outcomes which leads to a 2-dimensional one-hot token alphabet $\mathcal{A}$ with two different symbols. The Bayes predictors for all three tasks are analytically tractable and textbook examples (for the ‘Single coin’, the “Bayes” predictor is simply a constant probability).
@@
We evaluate both LSTMs and Decoder-only Transformers. To support all fine-tuning methods we always use an initial embedding, and a final unembedding layer. The embedding is a trainable linear projection from the 2D token space into a $128$-dimensional “embedding” space. The unembedding is a trainable linear projection from the outputs of the final network layer down to the 2D logits. Implementation Details. The LSTM has a single hidden layer of width $128$; the Transformer has a single multi-head attention layer with output dimensionality of $128$, $4$ attention heads, causal masking, SinCos positional encoding, a widening factor of $4$ for the MLP block, and layer normalization after query and key dense layers. LoRA fine-tuning (Lester et al., 2021) is only supported for the Transformer, where we apply LoRA to all dense matrices of the attention block (query, key, value, and final attention weights). To produce a prediction given the empty context, we pass an initial zero vector $P_{\theta}(x_{1}|\epsilon)=\pi_{\theta}(x_{1}|{\bf 0})$. This zero vector 0 is also prepended before any tunable prefix. When reporting the internal state of the LSTM, we use the cell state (hidden state gives qualitatively similar results), and for the Transformer we use the (causally masked) output of the attention block.
Figure 2: 2D PCA projection of Transformer’s (top) and LSTM’s (bottom) internal state (= activations), illustrating how differently tuned prefixes affect state and subsequent dynamics. Colored lines are sequences from target distribution (single coin with bias $0.2$), gray lines are from pretraining distribution (uniform random coins). The off-distribution nature of soft prefixes is particularly visible for the Real- and Soft-prefix for the Transformer. Fig. A7 shows that the vertical principal component corresponds to the step $n$, and the horizontal to the heads-to-tails ratio. See Fig. 1 for regret curves.
@@
Our main performance measure is the expected cumulative regret, which is the excess log loss compared to the ground-truth data generating probability. In Eq. 4 we defined the expected excess log loss relative to the best predictor that does not know $\tau$, a.k.a., the Bayesian regret. Similarly, we now define the excess log loss w.r.t. the data generator, that is, an oracle predictor that knows $\tau=\tau^{*}$:
@@
[…]
@@
We pretrain for $1000$ gradient steps (batch size $256$, sequence length $N_{\text{pre}}=100$, learning rate $0.001$, and gradient clipping if the norm is $\geq 1$). For tuning we use $1000$ steps (batch size of $256$, thus $K=256,000$, sequence length $N_{\text{tune}}=50$, learning rate of $5e^{-3}$, and gradient clipping if the norm is $\geq 1$). We show tuning loss curves (and their convergence) in the extended results in the appendix. We repeat tuning runs $10$ times per method with a different random seed (for sampling from the target distribution, and a different prefix initialization). Across repetitions and tuning methods, we always evaluate on the same set of $2048$ evaluation sequences. Results are reported as the median over repetitions with $25\%,75\%$ quantiles as ‘error bars’.
@@
We compare four different prefix-tuning methods and five different weight-tuning methods against a number of baselienes:
  • HardPT, SimplexPT, RealPT, SoftPT: prefix-tuning methods (see Section 3). To implement $\mathcal{S}=\Delta\mathcal{A}$ for SimplexPF, we pass the tunable prefix through a softmax. The prefix length ${L=6}$ in all main experiments, and we show $L=25$ in Appendix H.
@@
  • LoRAWT: Low-Rank Adaptation (Hu et al., 2021), where an additive tunable low rank matrix (rank=$4$ in our experiments) is added to all dense matrices of the attention block.
  • TargetBayes, PreBayes, PreBayesPT: exact Bayes predictors for the target distribution $\xi^{\text{Target}}$, the pretraining distribution $\xi^{\text{Pre}}$, and $\xi^{\text{Pre}}(\cdot|s^{\text{Target}}_{1:L})$, i.e., ‘PreBayes’ tuned to the target distribution (via exhaustive hard token search, $L=6$).
@@
(a) Top: Performance of different tuning methods, measured as excess log loss, i.e., regret (Eq. 10, lower is better). See bar plots for color legend. Top-right: Detailed Transformer results for last step within the tuning sequence length ${N_{\text{tune}}}$ and the last evaluation step ${N_{\text{eval}}}$. Right: Like above but for LSTM.
Figure 3: Models pretrained on sequences from coins with uniform random bias (length ${N_{\text{train}}=100}$) are fine-tuned to the target task of a mixture of two coins (tuning sequence length ${N_{\text{tune}}=50}$). No prefix-tuning method (with prefixes of length $6$) can achieve optimal performance on the target task (‘TargetBayes’ is optimal). Full weight-tuning, LoRA (on the Transformer) and two of the embedding tuning variants on the LSTM do reach optimality (even beyond the tuning length of $50$ steps). See Fig. A7 for a visualization of how different prefixes affect models’ internal dynamics. Regret curves for the LSTM, similar to Top left panel, are shown in Fig. A6.

### v2

@@ 1 Introduction
(a) Top: Performance of different tuning methods on Transformers, measured as excess log loss, i.e., regret (Eq.˜10, lower is better). See bar plots for color legend. Top-right: Detailed Transformer results for last step within the tuning sequence length $N_{\text{tune}}$} and the last evaluation step $N_{\text{eval}}$}. Right: Like above but for LSTM.
@@
Figure 1: Pretraining on sequences from coins with uniform random bias (length $N_{\text{train}}=100$}), then fine-tuning to the target task of a single coin with bias $0.2$ (tuning sequence length $N_{\text{tune}}=50$}). Plots show prediction performance on the target task for different prefix- and weight-tuning methods. For both Transformers and LSTMs Soft Prompting (‘SoftPT’) leads to optimal performance, showing that networks can be successfully prompted to behave Bayes-optimally on the target distribution (‘TargetBayes’). This holds up to the tuning sequence length ($50$), with only minor degradations up to $200$ steps. The corresponding soft prefixes of length=6=6 outperform even the best hard token prefixes of the same length (‘HardPF’). Most weight-tuning methods also perform very well. See Section˜4 for method details. Thick lines and bars show the median over $10$ tuning repetitions, thin lines individual repetitions, and shaded areas/bars show $25\%,75\%$ quantiles. See Fig.˜2 for a visualization of models’ internal dynamics. Regret curves for the LSTM, similar to top-left panel, are shown in Fig.˜A5.
In this paper, we investigate prefix-tuning, where a short prompt prefix is tuned to maximize subsequent prediction accuracy. We consider prefix search over hard tokens, and three soft prefix-tuning methods, including ‘Soft Prompting’, (Lester et al., 2021). The code to reproduce all our experiments is available at: https://github.com/google-deepmind/thunnini. Our main contributions are:
  • We discuss how prompting can be understood as steering a Bayesian sequential predictor via its pretrained in-context adaptation mechanism that arises from meta-training, see Section˜2.
  • We analyze theoretical conditions for the relationship between pretraining distribution and target task under which optimal prompting is and is not possible, see Section˜3.
  • We confirm these theoretical conditions empirically via a series of educational experiments, and show that in the negative case weight-based fine-tuning can succeed. See Section˜4.2.
  • We investigate mechanistic aspects of prompting LSTMs and Transformers, with experiments on both pretrained and untrained networks. In all cases, soft prefixes can be much more effective than any hard token sequence of equal length, particularly for Transformers, where soft prefixes can even cause untrained networks to behave as well-performing sequential predictors. See Section˜4.2.
@@
As with any conceptual principle, the Bayesian view cannot exhaustively describe all in-context learning phenomena at (frontier model) scale: limited data and model expressivity, suboptimal optimization, out-of-distribution generalization, and other mechanistic aspects will have additional impact. Having said that, there are a number of recent investigations at LLM scale supporting the plausibility of the Bayesian view, such as Chan et al. (2022a) who argue that distributional properties of natural language can easily give rise to a meta-learning setting. This paper discusses fundamental properties of prefix tuning, which we illustrate with educational experiments where the focus is on clarity and being able to compare quantitatively against a known and tractable Bayesian predictor. Accordingly, our datasets do not capture the full complexities of, e.g, large-scale language, vision, or robotics tasks. Similarly, our neural networks are small by modern standards, which means that our findings must be very cautiously extrapolated towards modern frontier model scale. Further rigorous and well-designed scientific studies are necessary to bridge the gap between our current work and modern large-scale ML practice, and we are optimistic that our fundamental results will inspire the design of such studies. The simplicity of our experimental setup allows to carefully control the data statistics, and compare neural predictors against exact Bayes-optimal predictors—both of which is not true for LLM-scale experiments. Additionally, it allows us to focus on fundamental aspects of prompting that arise even in idealized settings, and thus hold at any scale. We believe that this lays important fundamental groundwork, which will help future research to isolate and more effectively investigate additional, non-idealized, aspects of in-context learning and prompt tuning at frontier scale. We do not propose improved prompt tuning methods in this work since our aim is to study and understand existing ones.
@@ 2 Background: Memory-Based Meta-Learning
In this section we review how memory-based meta learning leads to Bayesian sequential predictors, whose hallmark feature is (most) rapid in-context adaptation. Under this view, the role of prompt tuning is to facilitate the inference process of the target task. While LLMs and frontier models are not explicitly meta-trained, their training process can be viewed as implicit meta-training, in which case the Bayesian view constitutes an important conceptual principle for understanding pretrained models and their in-context learning abilities.
@@
Let $\pi_{\theta}$ be a parametric sequential predictor, such as a neural network, which is a function with parameters $\theta$ that takes in an arbitrary-length sequence of $D$-dimensional float vectors and outputs a discrete distribution over the next (one-hot) token: $\pi_{\theta}:\{\mathbb{R}^{D}\}^{*}\rightarrow\Delta\mathcal{A}$ Typically, $D=|\mathcal{A}|$}, meaning that one-hot tokens can be fed directly into the predictor. The predictor’s conditional distribution over the next token given a context $x_{<n}$ is given by evaluating the predictor with the given context (i.e., performing a forward pass): $P_{\theta}(x_{n}|x_{<n})=\pi_{\theta}(x_{<n})$}. Keep in mind that a neural net accepts not only hard tokens as inputs, but sequences of arbitrary vectors in $\mathbb{R}^{D}$ (‘soft tokens’), and will produce a distribution over the next hard token in either case.
@@
which is zero iff $\pi_{\theta}=\xi$. Any predictor that fulfills this is Bayes-optimal for $\xi$, and Eq.˜1 and Eq.˜2 provide a recipe for constructing an explicit Bayesian predictor, which is a mixture over one predictor per task, weighted by the posterior probability over the task given the context so far: $P(\tau|x_{<n})\propto P(x_{<n}|\tau)P(\tau)$}. In many cases this recipe is analytically or computationally intractable. Memory-based meta-learning provides an alternative for obtaining an approximate Bayesian predictor simply through log loss minimization in a meta-learning loop:
@@
meaning the network’s prediction over the next token, given context $x_{<n}$, is (nearly) indistinguishable from an explicit Bayesian predictor. The meta-trained neural network thus implements a Bayes-optimal adaptive prediction algorithm via its activations only—without weight updates (in-context learning). Previous works have empirically verified that meta-trained LSTMs and Transformers can indeed reach Bayes-optimality through meta-training, e.g., Mikulik et al. (2020) for sequential prediction and decision-making (not covered in this paper, but the theory extends straightforwardly to loss functions other than log loss), Genewein et al. (2023) for piecewise stationary data sources (where models additionally have to infer task boundaries), and Grau-Moya et al. (2024) for variable-order Markov processes.
@@
Frontier models are implicitly meta-trained on samples from an unknown, rich and complex distribution of data generators (Chan et al., 2022a). The main concerns w.r.t. the applicability of the Bayesian viewpoint is that due to limited expressivity, limited data, suboptimal optimization, and off-distribution inputs, models may not converge to or operate in the Bayesian regime. Additionally, models may often operate in the generalization regime, while the theoretical guarantees only hold strictly under data drawn from the training meta-distribution. While we strongly recommend carefully investigating these issues, the theory tells us that as models get better and better, they will get closer and closer to the Bayesian ideal, making it an important fundamental computational mechanism to understand.
@@ 3 Prompt Optimization: prefix-tuning
We are given a neural sequential predictor $\pi_{\theta}$, that was pretrained [^1] Except for our experiments with untrained networks, where the parameters are at random initialization. via meta training over $\xi^{\text{Pre}}(x_{1:N})=\int P(x_{1:N}|\tau)P^{\text{Pre}}(\tau)d\tau$}. The goal is to adapt this predictor to a target distribution $\xi^{\text{Target}}(x_{1:N})=\int P(x_{1:N}|\tau)P^{\text{Target}}(\tau)d\tau$}. In prefix-tuning, the adaptation is performed by finding a (typically short) prefix sequence sequence $s_{1:L}\in\mathcal{S}^{L}$} of length $L$ that is prepended to the observations fed to the model. The “alphabet” $\mathcal{S}$ depends on the prefix-tuning method. In this paper we use:
@@
that is, we are optimizing for a single target task that had support under the pretraining distribution. In this case, there always exists a sequence of hard tokens that causes the Bayesian posterior to concentrate sufficiently (in the limit a delta, see Appendix˜E ) for optimal prediction after the prefix. For sufficiently large $L$,
@@
The prefix can be found by performing the minimization in Eq.˜7. Proof sketch: If the target distribution is a delta over one of the pretraining tasks (condition in Eq.˜8), then the argument that is being minimized in Eq.˜7 is the r.h.s. of Eq.˜9. The minimum is obtained when Eq.˜9 becomes an equality, which the case when the Bayesian (posterior) mixture $\xi^{\text{Pre}}(\cdot|s_{1:L})$} collapses to a single mixture component corresponding to $\tau^{\text{Target}}$. This also implies $D_{\text{KL}}(\xi^{\text{Target}}||\xi^{\text{Pre}}(\cdot|s_{1:L}))\approx 0$} which is only possible iff $\xi^{\text{Target}}\approx\xi^{\text{Pre}}(\cdot|s_{1:L})$.
If the condition in Eq.˜8 does not hold, optimal prompting may still be possible, but this strongly depends on the relationship between pretraining and target distribution and the model class. See Appendix˜E for an analysis of the Beta-Bernoulli case and beyond, including limits of prompting universal predictors. It is also possible to formulate two general theoretical negative cases:
@@
In the limit posteriors can often not remain (or become) multimodal. If the target distribution is multimodal, such as a finite target mixture over tasks, optimal prefix-tuning may not be possible, even if all mixture components have support under the pretraining distribution. For instance, if the prior $P^{\text{Pre}}(\tau)$} is log-concave and the likelihood function is log-concave, then the posterior is also log-concave, and thus unimodal. See Section˜4.2 for an empirical demonstration. More generally, if the posterior collapses to a Dirac delta in the limit (as it does for a Beta-Bernoulli model, and is very likely to do in general if prompts are typical sequences—see Appendix˜E), the only target task distributions that are optimally promptable are deltas over a single pretraining task.
@@
In general, for a prefix prompt to be optimal, the model’s internal state after consuming the prefix needs to be a sufficient statistic for the target distribution, without causing subsequent internal dynamics to diverge. Pretraining determines a model’s state-update function, and thus imposes strong constraints w.r.t. possible manipulations via hard token inputs. These constraints can be partly overcome by using soft prefixes instead of hard tokens. As our results show, these carefully tuned off-distribution inputs that lead to off-distribution internal states, can be used to very effectively steer pretrained, and even untrained neural predictors. The limits of this mechanistic aspect, outside the conceptual Bayesian theory, are currently unclear—it could be that soft prefixes can very flexibly “reprogram” a pretrained model to arbitrary target distributions. Empirically, we find this not to be the case: while Soft Prompting does consistently improve prediction performance on the target task, it is still bound by the theoretical limitations pointed out above and in detail in Appendix˜E: optimally adapting a predictor pretrained over uniform random coins to a mixture of two coins is not possible. How large the potential gains from Soft Prompting or other soft prefix-tuning methods can be is an empirical question. Finally, note that weight-based fine-tuning methods are able to modify the pretrained state-update and -readout mechanisms, which allows for more flexibility w.r.t. adapting a network to target distributions (see Section˜4.2 for empirical demonstrations).
@@ 4 Experiments on Coin-Flip Sequences
We conduct a series of experiments where a neural network is first meta-trained over a pretraining distribution (Section˜2) of coin flip sequences of length $N_{\text{pre}}=100$, and then prompt-tuned (Eq.˜7) or weight-tuned (mini-batch based log-loss minimization) to a target distribution of coin flip sequences of length $N_{\text{tune}}=50$. After tuning, we evaluate tuned models on $2048$ sequences of length $N_{\text{Eval}}=200$ from the target distribution. Choosing $N_{\text{Eval}}>N_{\text{tune}}$} also allows to study how the solutions of different tuning methods generalize beyond the tuning length. Across experiments we use three different data distributions, two neural architectures, and nine tuning methods, which we now describe.
@@ 4.1 Experimental setup
We use coin-flip sequences $P(x_{1:N}|\tau)=\text{Bernoulli}(\tau)$ with three different distributions $P(\tau)$ throughout our experiments. Random coins: $P{(\tau)=\text{Beta}(1,1)}$, leading to a uniform distribution over coin biases. This is our pretraining distribution. The exact Bayesian predictor in this case is the Laplace predictor. Single coin: A single coin with bias $0.2$. This target distribution fulfills the condition that makes optimal prompting possible in Eq.˜8. Two-coin mixture: A mixture of two coins, one with bias $0.2$ and one with bias $0.8$, with equal mixing weights of $1/2$ each. This target distribution violates the condition that theoretically allows for optimal prompting in Eq.˜8. All tasks have binary outcomes which leads to a 2-dimensional one-hot token alphabet $\mathcal{A}$ with two different symbols. The Bayes predictors for all three tasks are analytically tractable and textbook examples (for the ‘Single coin’, the “Bayes” predictor is simply a constant probability).
@@
We evaluate both LSTMs and Decoder-only Transformers. To support all fine-tuning methods we always use an initial embedding, and a final unembedding layer. The embedding is a trainable linear projection from the 2D token space into a $128$-dimensional “embedding” space (results for $4$-dimensional embeddings are shown in Appendix˜I). The unembedding is a trainable linear projection from the outputs of the final network layer down to the 2D logits. Implementation Details. The LSTM has a single hidden layer of width $128$; the Transformer has a single multi-head attention layer with output dimensionality of $128$, 44 attention heads, causal masking, SinCos positional encoding, a widening factor of $4$ for the MLP block, and layer normalization after query and key dense layers. Results for larger networks are shown in Appendix˜J. LoRA fine-tuning (Lester et al., 2021) is only supported for the Transformer, where we apply LoRA to all dense matrices of the attention block (query, key, value, and final attention weights). To produce a prediction given the empty context, we pass an initial zero vector $P_{\theta}(x_{1}|\epsilon)=\pi_{\theta}(x_{1}|{\bf 0})$. This zero vector 0 is also prepended before any tunable prefix. When reporting the internal state of the LSTM, we use the cell state (hidden state gives qualitatively similar results), and for the Transformer we use the causally masked output of the attention block.
@@
Our main performance measure is the expected cumulative regret, which is the excess log loss compared to the ground-truth data generating probability. In Eq.˜4 we defined the expected excess log loss relative to the best predictor that does not know $\tau$, a.k.a., the Bayesian regret. Similarly, we now define the excess log loss w.r.t. the data generator, that is, an oracle predictor that knows $\tau=\tau^{*}$:
@@
We show regret curves from $N=0$ up to $N=N_{\text{eval}}-1=199$ steps [^2] The offset of $1$ is because passing $x_{1:n-1}$ through the network produces a prediction $x_{n}$ for which we need ground-truth data at step $n$ to compute the regret or log loss (or gradients).. For prefix-tuning methods $\tilde{\theta}$ refers to the pretrained weights (or randomly initialized weights in our experiments on untrained networks) and $s_{1:L}$ is the tuned prefix. For weight-tuning methods $\tilde{\theta}$ refers to the tuned weights and the prefix is empty ($L=0$). We (Monte-Carlo) estimate the regret with $2048$ sequences sampled from the target data generator (from which we also get the ground-truth generating probabilities).
@@
We pretrain for $1000$ gradient steps (batch size $256$, sequence length $N_{\text{pre}}=100$, learning rate $0.001$, and gradient clipping if the norm is $\geq 1$). For tuning we use $1000$ steps (batch size of $256$, thus $K=256,000$, sequence length $N_{\text{tune}}=50$, learning rate of $5e^{-3}$, and gradient clipping if the norm is $\geq 1$). We show tuning loss curves (and their convergence) in the extended results in the appendix. We repeat tuning runs $10$ times per method with a different random seed (for sampling from the target distribution, and a different prefix initialization). Across repetitions and tuning methods, we always evaluate on the same set of $2048$ evaluation sequences. Results are reported as the median over repetitions with $25\%,75\%$ quantiles as “error bars”.
@@
We compare four different prefix-tuning methods and five different weight-tuning methods against a number of baselines:
  • HardPT, SimplexPT, RealPT, SoftPT: prefix-tuning methods (see Section˜3). To implement $\mathcal{S}=\Delta\mathcal{A}$ for SimplexPF, we pass the tunable prefix through a softmax. The prefix length L=6{L=6} in all main experiments, and we show $L=25$ in Appendix˜H.
@@
  • LoRAWT: Low-Rank Adaptation (Hu et al., 2021), where an additive tunable low rank matrix (rank=$4$ in our experiments) is added to all dense matrices of the attention block. All other weights remain frozen.
  • TargetBayes, PreBayes, PreBayesPT: exact Bayes predictors for the target distribution $\xi^{\text{Target}}$, the pretraining distribution $\xi^{\text{Pre}}$, and $\xi^{\text{Pre}}(\cdot|s^{\text{Target}}_{1:L})$, i.e., ‘PreBayes’ prefix-tuned to the target distribution via exhaustive hard token search, $L=6$.
@@
Figure 2: 2D PCA projection of Transformer’s (top) and LSTM’s (bottom) internal state (= activations), illustrating how differently tuned prefixes affect state and subsequent dynamics. Fig.˜A7 shows that the vertical principal component corresponds to the step $n$, and the horizontal to the heads-to-tails ratio. Colored lines are sequences from the target distribution (single coin with bias $0.2$), gray lines are from the pretraining distribution (uniform random). The off-distribution nature of soft prefixes is particularly visible for the Real- and Soft-prefix for the Transformer. See Fig.˜1 for regret curves.
(a) Top: Performance of different tuning methods, measured as excess log loss, i.e., regret (Eq.˜10, lower is better). See bar plots for color legend. Top-right: Detailed Transformer results for last step within the tuning sequence length $N_{\text{tune}}$} and the last evaluation step $N_{\text{eval}}$}. Right: Like above but for LSTM.
Figure 3: Models pretrained on sequences from coins with uniform random bias (length $N_{\text{train}}=100$}) are fine-tuned to the target task of a mixture of two coins (tuning sequence length $N_{\text{tune}}=50$}). No prefix-tuning method (with prefixes of length 66) can achieve optimal performance on the target task (‘TargetBayes’ is optimal). Full weight-tuning, LoRA (on the Transformer) and two of the embedding tuning variants on the LSTM do reach optimality (even beyond the tuning length of $50$ steps). See Fig.˜A7 for a visualization of how different prefixes affect models’ internal dynamics. Regret curves for the LSTM, similar to Top left panel, are shown in Fig.˜A6.